# D4C glove-train: solving the RPM and Bongard-logo problem by distributing and Circumscribing concepts

### 17 Mar 2024 v3

@@ Abstract
This paper achieves significant progress in the field of abstract reasoning, particularly in addressing Raven’s Progressive Matrices (RPM) and Bongard-Logo problems. We propose the D2C approach, which redefines conceptual boundaries in these domains and bridges the gap between high-level concepts and their low-dimensional representations. Based on this, we further introduce the D3C method that handles Bongard-Logo problems and significantly improves reasoning accuracy by estimating the distribution of image representations and measuring their Sinkhorn distance. To enhance computational efficiency, we introduce the D3C-cos variant, which provides an efficient and accurate solution for RPM problems by constraining distribution distances. Additionally, we present Lico-Net, a network that combines D3C and D3C-cos to achieve state-of-the-art performance in both problem-solving and interpretability. Finally, we extend our approach to D4C, employing adversarial strategies to further refine conceptual boundaries and demonstrate notable improvements for both RPM and Bongard-Logo problems. Overall, our contributions offer a new perspective and practical solutions to the field of abstract reasoning.
@@ III Methodology
This paper designs a new baseline model named Lico-Net for the RPM problem and achieves considerable performance on this problem. In addition, this paper extends the existing Bongard-Logo baseline in terms of model size and observes its limitations. Inspired by the limitations of the Bongard-Logo baseline, this paper designs the D2C method, which can be widely applied to RPM problems and Bongard-Logo, leading to improvements in model performance. This paper also designs the D3C method, which adheres to a core idea: representing complex, highly abstract human concepts with distributions rather than vectors is more appropriate. Finally, this paper designs the D4C method, which aims at the essence of abstract reasoning problems and contributes to RPM problems and Bongard-Logo problems.
@@ IV-A A novel baseline, Lico-Net, for RPM problem
Aligned with the current consensus among RPM solvers, which emphasizes the need for multi-scale or multi-viewpoint feature extraction from RPM images[28, 23, 24], Lico-Net incorporates the Vision Transformer (ViT) as a perceptron for feature extraction, while preserving the entire output vector, attention results, as multi-view point features. In this paper, we denote the number of viewpoints as $L$, a practice similar to that employed in Savirt[24], RS-Tran[28] and Triple-CFN. Lico-Net subsequently processes each viewpoint of extracted feature equally.
@@
For the isometric viewpoint of an completed $3\times 3$ progressive matrix, we employ an MLP (Multi-layers perceptron) with a bottleneck structure to extract information from all minimal reasoning units (three images within a row for RAVEN and three images within a row or column for PGM) and obtain unit vectors. Specifically, we concatenate the representations of all images within a minimal reasoning unit in their inherent order and input them into the MLP. Compared to other structures, the MLP more directly preserves the order of the input representations and its structure is sufficiently simple. Due to the design of RPM problems, typically multiple minimal reasoning units must be combined to arrive at the image progression pattern of a RPM matrix. For instance, a RAVEN case requires at least two row unites to discern the case’s pattern, while PGM necessitates at least two row unites and two column unites[39]. It’s worth noting that determining the RPM matrix’s progression pattern doesn’t necessitate all minimal reasoning units; therefore, numerous combination methods of minimal reasoning units exist that are sufficient to identify the pattern. We refer to the combination requiring the least unit vectors as the minimal concept group.
@@
Given that RPM problems might be constrained by multiple independent and decouple patterns, we embed $N$ optimizable vectors within each minimal concept group (in this paper, we embed two optimizable vectors, with $N$ being a hyperparameter that can be freely set) and input them into a standard Transformer-Encoder as tokens to compute $N$ progression pattern vectors for the progressive matrix in the case. Assuming there are $M$ minimal concept groups, we have thus obtained $N\times M$ progression pattern vectors under one viewpoint. We denote the progressive pattern vector from a single perspective as $\{P_{nm}|\,n\in[1,N],m\in[1,M]\}$ These vectors have additional purposes to fulfill. This paper independently calculates the scores of each option in the problem under $N$ distinct pattern vectors.
@@
The $M$ minimal concept groups can be divided into two categories: those involving options and those not involving options. In this paper, we select the pattern vectors calculated from the concept groups not involving options as queries, and the pattern vectors involving options as keys and values. We then compute the multi-head cross-attention results between them and use a new MLP to score the attention outputs. Since we intentionally introduced $N$ patterns, we obtain $N$ scores, which are summed to yield the final score for one viewpoint. Finally, we calculate the overall score of one option by averaging the scores from all viewpoints. Aligned with the answer logic, we employ the Cross-Entropy loss function to constrain these scores for optimizing Lico-Net. It’s worth noting that we treat each dimension of the unit vector as a distribution and use covariance matrices as additional loss function term to control the correlation between these distributions, aiming to encourage the unit vectors to contain more information. The loss function $\ell_{cov}$ attached to a batch of the minimal reasoning unit vectors $\{U_{ij}|\,i\in[1,b],j\in[1,a]\}$ can be expressed as:
@@ IV-B The baseline for Bongard-Logo problem
V Dynamic Delineation of the Concept (D2C)
@@
The determination of image attributes is significantly influenced by the specific contextual concepts associated with each individual case. Therefore, this paper aims to develop a method that effectively decouples the concepts across different cases. To accomplish this goal, we redefine the concepts within the Bongard-Logo database, representing them as conditional probability distributions $p_{i}(x|\,Y_{i})$ and $q_{i}(x|\,Y_{i})$. Here, $i$ ranges from $1$ to $n$, representing each index of Bongard-Logo cases. Additionally, we introduce the notation $\{Y^{\prime}_{i}|\,i\in[1,n]\}$ to denote the set of reorganized concepts. It can be seen that the number of reorganized concepts corresponds to the number of Bongard-Logo cases. Through this concept reorganization, we strive to achieve mutually exclusive expressions of concepts among all Bongard-Logo cases. In other words, each case is uniquely categorized based on its own distinct set of concepts, with corresponding unparalleled classification rules. By establishing a one-to-one relationship between image styles and these redefined concepts, we hope to enable the network to seamlessly infer concepts from the pixel arrangements of images.
@@
In pursuit of fostering concept antagonism between differing problem domains, the present study has revised the data sampling techniques employed in mini-batch training methodologies. Furthermore, an additional loss function term has been designed to facilitate the reassignment of concepts within the network architecture. This method, denoted as “D2C”(Dynamic Delineation of the Concept), strives to cultivate a state of concept contrast amidst diverse problem sets.
@@
V-A D2C phase 1: The concept delineation method based on a novel dynamic mini-batch sampling technique
Through a meticulous selection process, this paper identifies mutually exclusive pairs of Bongard-logo problems for training purposes and establishes concept antagonism within this triplet using the InfoNCE loss function.
@@ V-A D2C phase 1: The concept delineation method based on a novel dynamic mini-batch sampling technique
Specifically, we encode a batch of Bongard-logo cases using ResNet50 and obtain a representation set $\{Z_{ij}|\,i\in[1,b],j\in[1,14]\}$, where $i$ denotes the index of the Bongard-logo case within a batch, the batch size is $b$, and $j$ represents the sample index within a single Bongard-logo case, with $j\in[1,7]$ being the primary samples and $j\in[8,14]$ being the auxiliary samples. From this batch of cases $\{Z_{ij}\}$, we deliberately choose two batches of cases $\{Z_{tj}|\,j\in[1,14]\}$ and $\{Z_{\tilde{t}j}|\,j\in[1,14]\}$, ensuring that $t$, $\tilde{t}$, and $i$ are all distinct. This process can be viewed as ancestral sampling. The extracted primary samples $\{Z_{tj}|\,j\in[1,7]\}$ and $\{Z_{\tilde{t}j}|\,j\in[1,7]\}$ are then incorporated into $\{Z_{ij}\}$ as auxiliary samples, yielding an expanded set $\{Z_{i\tilde{j}}\}$ where $\tilde{j}\in[1,28]$. Subsequently, the InfoNCE loss function is computed for these newly formed sample groups $\{Z_{i\tilde{j}}|\,i\in[1,b],\tilde{j}\in[1,28]\}$. The feedforward process, as illustrated in the accompanying diagram, involves a strategic misalignment of Bongard-logo problem batches to facilitate ancestral sampling during implementation. This process can be expressed as figure 6.
@@
V-B D2C phase 2: The concept delineation method based on a new loss function term

### 25 Mar 2024 v4

# D4C Glove-train: Solving the RPM and Bongard-logo Problem by Circumscribing and Building Distribution for Concepts

@@ Abstract
This paper achieves noteworthy progress in the realm of abstract reasoning, particularly in addressing Raven’s Progressive Matrices (RPM) and Bongard-Logo challenges. Initially, we introduce Lico-Net, a novel baseline model that resolves RPM problems with remarkable accuracy. Leveraging this foundation, we advance with the D3C approach, which advocates representing the underlying concepts in abstract reasoning problems through distributions. This perspective enhances the performance of both Lico-Net and a baseline model excelling in Bongard-Logo tasks. To bolster the computational efficiency of D3C, we present the D3C-cos variant, offering a streamlined yet precise solution. Furthermore, we propose the D2C method, redefining conceptual boundaries within these domains and bridging the divide between high-level abstractions and their lower-dimensional counterparts. Finally, we extend our methodology to D4C, employing adversarial techniques to refine conceptual boundaries further and demonstrate substantial improvements in both RPM and Bongard-Logo challenges. Overall, our contributions present a fresh outlook and practical advancements in the field of abstract reasoning.
@@ III Methodology
This paper presents Lico-Net, a novel baseline model specifically designed for the RPM problem, which achieves remarkable performance in addressing this cognitive challenge. Subsequently, we introduce a fundamental argument that representing abstract and complex concepts as distributions, rather than vectors, is a more suitable approach. Based on this insight, we develop the D3C method. This technique not only enhances Lico-Net’s performance and interpretability on RPM tasks but also elevates the reasoning accuracy of Triple-CFN, a notable deep model that has excelled in Bongard-Logo problems. Recognizing the sources of reasoning difficulties in Bongard-Logo problem, we devise the D2C method. This method is broadly applicable to both RPM and Bongard-Logo challenges, demonstrating improved performance over several baseline models in these domains. Finally, we introduce the D4C method, which targets the essential aspects of abstract reasoning. This approach contributes significantly to advancing our understanding of RPM and Bongard-Logo problems, paving the way for future research in abstract reasoning.
@@ IV-A1 Perceptron module of Lico-Net
Aligned with the current consensus among RPM solvers, which emphasizes the need for multi-scale or multi-viewpoint feature extraction from RPM images[28, 23, 24], Lico-Net incorporates the Vision Transformer (ViT) as a perceptron for feature extraction, while preserving the entire output vector, attention results, as multi-view point features. In this paper, we denote the number of viewpoints as $L$, a practice similar to that employed in Savirt[24], RS-Tran[28] and Triple-CFN. Lico-Net subsequently processes each viewpoint of extracted feature equally. The perceptron of Lico-Net is illustrated in figure 4
Figure 4: Perceptron of Lico-Net
IV-A2 Reasoning module of Lico-Net
The Reasoning module of Lico-Net processes each view point feature generated by the perception module equally and in parallel, resulting in $L$ sets of inference outcomes. Subsequently, the mean value of these inference results is utilized as the global inference outcome for Lico-Net. The detailed process is described as follows.
@@ IV-A2 Reasoning module of Lico-Net
For the isometric viewpoint of an completed $3\times 3$ progressive matrix, we employ an MLP (Multi-layers perceptron) with a bottleneck structure to extract information from all minimal reasoning units (three images within a row for RAVEN and three images within a row or column for PGM) and obtain unit vectors. Specifically, we concatenate the representations of all images within a minimal reasoning unit in their inherent order and input them into the MLP. Compared to other structures, the MLP more directly preserves the order of the input representations and its structure is sufficiently simple. The extraction process of unit vector is illustrated in figure 5.
Figure 5: The extraction process of unit vector
Due to the design of RPM problems, typically multiple minimal reasoning units must be combined to arrive at the image progression pattern of a RPM matrix. For instance, a RAVEN case requires at least two row unites to discern the case’s pattern, while PGM necessitates at least two row unites and two column unites[39]. It’s worth noting that determining the RPM matrix’s progression pattern doesn’t necessitate all minimal reasoning units; therefore, numerous combination methods of minimal reasoning units exist that are sufficient to identify the pattern. We refer to the combination requiring the least unit vectors as the minimal concept group.
@@
Given that RPM problems might be constrained by multiple independent and decouple patterns, we embed $N$ optimizable vectors within each minimal concept group (in this paper, we embed two optimizable vectors, with $N$ being a hyperparameter that can be freely set) and input them into a standard Transformer-Encoder as tokens to compute $N$ progression pattern vectors for the progressive matrix in the case. Assuming there are $M$ minimal concept groups, we have thus obtained $N\times M$ progression pattern vectors under one viewpoint. We denote the progressive pattern vector from a single perspective as $\{P_{nm}|\,n\in[1,N],m\in[1,M]\}$ These vectors have additional purposes to fulfill. This calculation process can be expressed in figure 6.
Figure 6: The extraction process of pattern vector
This paper independently calculates the scores of each option in the problem under $N$ distinct pattern vectors.
@@
The $M$ minimal concept groups can be divided into two categories: those involving options and those not involving options. In this paper, we select the pattern vectors calculated from the concept groups not involving options as queries, and the pattern vectors involving options as keys and values. We then compute the multi-head cross-attention results between them and use a new MLP to score the attention outputs. Since we intentionally introduced $N$ patterns, we obtain $N$ scores, which are summed to yield the final score for one viewpoint. This process is illustrated in figure 7.
Figure 7: The calculation process of final score for one viewpoint
Finally, we calculate the overall score of one option by averaging the scores from all viewpoints. Aligned with the answer logic, we employ the Cross-Entropy loss function to constrain these scores for optimizing Lico-Net. It’s worth noting that we treat each dimension of the unit vector as sample results from a distribution and use covariance matrices as additional loss function term to control the correlation between these distributions, aiming to encourage the unit vectors to contain more information. The loss function $\ell_{cov}$ attached to a batch of the minimal reasoning unit vectors $\{U_{ij}|\,i\in[1,b],j\in[1,a]\}$ can be expressed as:
@@ IV-B The baseline for Bongard-Logo problem
V Distribution Distance-based Discriminator of Case (D3C)
V-A D3C for Bongard-Logo
The ideal approach to address the Bongard-Logo problem involves training an encoder capable of encoding images from a Bongard-Logo case into representations, denoted as $z$. In this context, the representations of positive instances (samples in primary group) belong to the distribution $p(z)$, while the representations of negative instances (samples in primary group) belong to $q(z)$. Importantly, there should be minimal overlap between $p(z)$ and $q(z)$ in terms of cross-measure. Therefore, this paper proposes the D3C method to enhance the performance of encoders in the form of $f(z|x)$ on Bongard-Logo problems. Notably, this brings improvements for models such as Triple-CFN and ResNet50.
Specifically, this paper hypothesizes that the representation of the primary group samples in a Bongard-Logo case follows a Gaussian distribution, while the representation of the auxiliary group samples follows a mixture of Gaussian distributions. We utilize an arbitrary backbone network $f(z|\,x)$ to approximate the distribution of image representations within Bongard-Logo cases. To achieve this, we modify the network to output parameters $\mu$ and $\sigma$ instead of a single vector, effectively converting $f(z|\,x)$ into $f(\mu,\sigma|\,x)$. The encoded vector is then divided into two equal-length segments, which we interpret as the mean and the logarithm of the variance, respectively. This approach normalizes the output vectors into Gaussian distributions and facilitates the distribution of image representations in Bongard-Logo cases. Our method bears resemblance to the Variational Autoencoder (VAE) framework. We argue that using distributions, rather than vectors, to describe image representations offers greater flexibility and accuracy in capturing the underlying complexity and uncertainty of abstract reasoning problems.
With the estimated distributions of representations for each image in the Bongard-Logo cases, we compute the Sinkhorn distance between these distributions using reparameterization technique. Following the logical structure of each Bongard-Logo case, we impose constraints on these distances for the backbone network $f(\mu,\sigma|\,x)$, aiming to train an effective discriminator. Specifically, we optimize the network by utilizing a binary cross-entropy loss function to enforce constraints on the exponential function of the negative Sinkhorn distance with base $e$ ($e^{-D}\in[0,1]$, where $D\in[0,+\infty]$ is the values of Sinkhorn distance), aligning with the case’s underlying logic mentioned in DC. This process can be expressed as figure 10. The D3C Method, applied to a batch of Bongard-Logo case distributions encoded by the backbone network $f(\mu,\sigma|\,x)$, can be expressed as follows:
${\ell_{\text{D3C}}}(\{z_{ijs}|\,i\in[1,b],j\in[1,14],s\in[1,S]\})$
Where $b$ is the batchsize of training. $S$ is set to 20, and $S$ represents the sampling frequency of the reparameterization technique.
Figure 10: Feedforward process of D3C
The design of D3C allows for its application to any image encoder that can be abstracted as $f(z|x)$. This requirement is met by a wide range of encoders, including those from the ResNet family, ViT, and specialized ones like Triple-CFN, which is particularly suitable for Bongard-Logo problems.
V-B D3C on Lico-Net for RPM
Notably, in contrast to the Bongard-Logo database, both RAVEN and PGM possess well-defined and comprehensive Meta data in the form of (auxiliary) rule labels. These rule descriptions facilitate the construction of a rational concept space, with boundary concepts within the rule labels. This paper introduces the rule labels from the RPM database into the Lico-Net, aiming to enhance the reasoning accuracy and human-perspective rule interpretability of Lico-Net. This process will be implemented by the D3C method.
Within the frameworks of RAVEN and PGM, there is a many-to-one mapping between image progression patterns and their corresponding rules. For instance, both single-step and double-step increments, as well as single-step and double-step decays in image attributes, are all categorized under the ‘progressive’ rule. Therefore, this paper proposes the concept of rules from the RPM database and represent them as distributions while also vectorizing the image progression patterns. We can leverage this approach to naturally constrain the image progression patterns within their corresponding rule distributions, thereby optimizing the deep model’s learning of rules and clarifying the boundaries among various concepts within the data.
In practice for D3C, we utilized a standard Transformer-Encoder to convert the textual descriptions of Meta data into concept vectors $\{{r}_{k}|\,k\in[1,K]\}$, where $K$ represents the count of distinct rule description in Meta data. Specifically, the initial portion of the feature vector is designated as the mean $\{\mu_{k}|\,k\in[1,K]\}$, whereas the latter half signifies the logarithm of variance $\{\log({\sigma_{k}}^{2})|\,k\in[1,K]\}$, effectively distributing the rule descriptions and yielding a set of distributions $\{{\mu}_{k},{\sigma^{2}}_{k}|\,k\in[1,K]\}$. Lico-Net has inherently achieved the vectorization of image progression patterns, yielding vectors termed as progression pattern vectors. Consequently, we integrated a constraint term into Lico-Net, dictating that vectors representing progression patterns should exhibit a high likelihood of adhering to the distribution outlined by their corresponding textual rules, and conversely, demonstrate a low probability of deviation.
A specific RPM case is encoded into $N\times M$ progression pattern vectors $\{P_{nm}|\,n\in[1,N],m\in[1,M]\}$ from $L$ viewpoints by Lico-Net as aforementioned. To align Lico-Net with the D3C framework, the number of encoding patterns $N$ in Lico-Net is set according to the instance of the RPM problem, depending on the number of decoupled concepts present in the Meta data of that instance. For example, in PGM, there are two such concepts: shape pattern and line pattern, and $N$ is set to 2. The special setting of $N$ in Lico-Net can be viewed as $N$ slots, which can be aligned with the $N$ decoupled concepts described in the Meta data, respectively. By synthesizing all perspectives—specifically, by calculating the mean of the progression pattern vectors $\{P_{nm}|\,n\in[1,N],m\in[1,M]\}$ across all $L$ viewpoints—Lico-Net encodes $N\times M$ vectors, denoted as $\{\overline{P}_{nm}|\,n\in[1,N],m\in[1,M]\}$, representing the full-viewpoint progression patterns. This process can be expressed as:
$\{\overline{P}_{nm}|\,n\in[1,N],m\in[1,M]\}$
$=\frac{1}{L}\sum_{l=1}^{L}\{P_{nm}|\,n\in[1,N],m\in[1,M]\}_{l}$ (7)
The $l$ represents the index of viewpoints. The $\{\overline{P}_{nm}|\,n\in[1,N],m\in[1,M]\}$ can be treated as $N$ sets of $M$ sampling results. This enables the optimization of distribution distances based on the Sinkhorn distance, thereby enhancing the overall performance through D3C framework. Apparently, this paper utilizes the Sinkhorn distance to constrain the distance between $M$ progression pattern vectors and their corresponding distributions’ sampling results. The $N$ decoupling concept presented in Meta data is employed to concurrently impose constraints on the $N$ sets of $M$ progression pattern vectors, and the process of aligning $\{\overline{P}_{nm}\}$ with the distributional concepts $\{{\mu}_{k},{\sigma^{2}}_{k}|\,k\in[1,K]\}$ can be illustrated in figure 11.
Figure 11: Feedforward process of D3C
The calculate porcess can be exprecessed as:
The $\overline{P}_{nm}$ represents the full-viewpoint progressive pattern vectors encoded by Lico-Net, and $\{r_{ks}\}$ represents the $S$ sampling results from the distribution $\{\mu_{k},{\sigma^{2}}_{k}\}$ of the rule description. It is evident that the specific value of $\tilde{k}$ is contingent upon the description of Meta data corresponding to the particular case of the RPM problem. symbol “$\tilde{k}|\,n$” represents that the specific value of $\tilde{k}$ is depended on $n$. This dependency realizes the fact that each slots in Lico-Net are aligned with their corresponding concepts. The $\text{Sinkhorn}(\cdot,\cdot)$ represents the Sinkhorn distance operator. Other logic of symbol in function (V-B) has mentioned. It becomes evident that we are employing the Binary Cross Entropy (BCE) loss function to impose constraints on the exponential negative Sinkhorn distance result. The feedforward process is illustrated in the figure 12.
Figure 12: Feedforward process of Lico-Net aligned with D3C
V-C A low-cost loss function for the D3C, the Hermes for the Zeus
In our study, we opted for cosine similarity, considering normalized vectors $u$ and $v$, where $||v||$ and $||u||$ are equal to 1, and Lipschitz constant $L=||v||=1$. It is reasonable to choose the mean of cosine similarity between sampling vectors as the distance metric function instead of Sinkhorn distance. The calculate process of new method applied on Bongard-Logo can be expressed as:
${\ell_{\text{D3C-cos}}}(\{z_{ijs}|\,i\in[1,b],j\in[1,14],s\in[1,S]\})$
The $t$ is set to $10^{-6}$ and $S$ is set to $5$ in function (V-C) and (V-C). The logic of symbols in function (V-C) and (V-C) has mentioned. It is evident that, compared to function (V-A) and (V-B), the demand for sampling frequency $S$ in D3C-cos has been significantly reduced.
The success of such simplification can be attributed, in part, to our modeling of the distribution of pattern descriptions as a Gaussian distribution.
VI Dynamic Delineation of the Concept (D2C)
In this section, the paper initially endeavors to introduce novel constraints for the Bongard-Logo discriminators, aiming to enhance their performance. Motivated by these constraints, we incorporate innovative designs into the proposed Lico-Net architecture.
VI-A D2C on Bongard-Logo problem
In the Bongard-Logo problem, the determination of image attributes is significantly influenced by specific contextual concepts associated with each individual case. Therefore, this study aims to develop a methodology that effectively decouples the concepts across different cases within the Bongard-Logo problem.
@@ VI-A D2C on Bongard-Logo problem
To achieve this objective, we have redefined the concepts within the Bongard-Logo database, representing them as conditional probability distributions $p_{i}(x|\,Y_{i})$ and $q_{i}(x|\,Y_{i})$. Here, $i$ ranges from $1$ to $n$, denoting each index of the Bongard-Logo cases. Additionally, we introduce the notation $\{Y^{\prime}_{i}|\,i\in[1,n]\}$ to represent the set of reorganized concepts. It is evident that the number of reorganized concepts corresponds to the number of Bongard-Logo cases. Through this concept reorganization, we strive for mutually exclusive expressions of concepts across all Bongard-Logo cases. In other words, each case is uniquely categorized based on its distinct set of concepts. By establishing a one-to-one correspondence between image styles and these redefined concepts, we aim to enable the network to seamlessly infer concepts from the pixel arrangements of images. This, in turn, aims to mitigate the impact of multiple attribute identifications on the solution of the Bongard-Logo problem.
@@
To foster concept antagonism between different problem domains, we have revised the data sampling techniques employed in mini-batch training methodologies. Furthermore, an additional loss function term has been devised to facilitate the reassignment of concepts within the network architecture. This approach, termed “D2C” (Dynamic Delineation of Concepts), endeavors to cultivate a state of concept contrast amidst diverse problem sets.
@@
VI-A1 D2C phase 1: The concept delineation method based on a novel dynamic mini-batch sampling technique
By augmenting the traditional sampling methods used for mini-batch training with an additional ancestral sampling procedure, this study identifies mutually exclusive pairs of Bongard-logo problems for training purposes. This constitutes a hot-swappable approach for training the Bongard-Logo solver, which is designated as phase 1 of the D2C method in this paper.
@@ VI-A1 D2C phase 1: The concept delineation method based on a novel dynamic mini-batch sampling technique
Specifically, we denote a batch of Bongard-logo cases as $\{x_{ij}|\,i\in[1,b],j\in[1,14]\}$, where $i$ denotes the index of the Bongard-logo case within a batch, the batch size is $b$, and $j$ represents the sample index within a single Bongard-logo case, with $j\in[1,7]$ being the primary samples and $j\in[8,14]$ being the auxiliary samples. From this batch of cases $\{x_{ij}\}$, we deliberately choose two batches of cases $\{x_{tj}|\,j\in[1,14]\}$ and $\{x_{\tilde{t}j}|\,j\in[1,14]\}$, ensuring that $t$, $\tilde{t}$, and $i$ are all distinct. This process can be viewed as ancestral sampling. The extracted primary samples $\{x_{tj}|\,j\in[1,7]\}$ and $\{x_{\tilde{t}j}|\,j\in[1,7]\}$ are then incorporated into $\{x_{ij}\}$ as auxiliary samples, yielding an expanded set $\{x_{i\tilde{j}}\}$ where $\tilde{j}\in[1,28]$. Subsequently, The new sample groups $\{x_{i\tilde{j}}\}$, which oppose the concepts from different cases compared to their original form, can be fed into any discriminator used to solve Bongard-Logo problems, such as Triple-CFN[32] and PMoC[31]. The feedforward process, as illustrated in the accompanying diagram, involves a strategic misalignment of Bongard-logo problem batches to facilitate ancestral sampling during implementation. The alignment of the D2C phase 1 process with the baseline in DC is illustrated in figure 13.
@@
In summary, the phase 1 of the D2C approach embodies a straightforward yet effective concept: it introduces positive examples from other cases as negative instances in the current case. This strategy sends a clear and additional signal to Bongard-Logo discriminators that distinct problem cases should rely on differing concepts. As a result, the first phase of D2C is widely applicable to various Bongard-Logo discriminators, including Triple-CFN and PMoC.
VI-A2 D2C phase 2: The concept delineation method based on a new loss function term